Multiple Choice Question 132 O $150000. O $135000. O $141000. O $120000 Question Aftempts 0 of i ed SAVE FOR LATER 0
Answer #1
Particulars Amount
Opening Work in Progress 138000
Add: Total Manufacturing Costs 577000
Less: Closing Work in Progress XX (To find out)
Cost of Goods Manufactured 580000
Opening WIP+Total Manufacturing Cost-Closing WIP=Cost of Goods Manufactured
138000+577000-Closing WIP=580000
715000-Closing WIP=580000
Closing WIP=715000-580000
Closing WIP=715000-580000
Closing WIP=135000
Answer is $135000/-